OpenStudy (anonymous):

Integrate the function\[\int\limits_{}^{}\frac{ 1 }{ t ^{2}\sqrt{2-t ^{2}} }\]

OpenStudy (anonymous):

Have you learned integration by parts yet?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

is that what i should use?

OpenStudy (anonymous):

I think so, I haven't worked it out yet... but we can try that together if you like. So in this example what would you choose for your "u" and which for your "dv"?

OpenStudy (anonymous):

u would be \[\frac{ 1 }{ t ^{2} }\] dv would be \[\frac{ 1 }{ \sqrt{2-t ^{2}} }\]

OpenStudy (anonymous):

keep in mind that we need to be able to integrate our "dv".

OpenStudy (anonymous):

so i suppose we want to switch those

OpenStudy (anonymous):

\[\frac{ 1 }{ \sqrt{2-t ^{2}} } \] is not easy to integrate.

OpenStudy (anonymous):

right

OpenStudy (anonymous):

so lets switch them. If dv=t^-2 then what is v? If u=the rest what is du?

OpenStudy (turingtest):

for the record, I got the answer using a trig substitution it may be doable with integration by parts, too

OpenStudy (anonymous):

trig subtstitution? please explain

OpenStudy (turingtest):

have you done tirg subs yet?

OpenStudy (turingtest):

trig*

OpenStudy (anonymous):

i've done trigonometric integrals

OpenStudy (anonymous):

oh wait yes i have

OpenStudy (turingtest):

\[\int{dt\over t^2\sqrt{2-t^2}}=\frac1{\sqrt2}\int{dt\over t^2\sqrt{1-(\frac t{\sqrt2})^2}}\]now let\[\frac t{\sqrt2}=\sin\theta\implies dt=\sqrt2\cos\theta d\theta\]

OpenStudy (anonymous):

The integration by parts becomes pretty messy stick with Turing.

OpenStudy (turingtest):

it usually does when you have both variable in the denom...

OpenStudy (turingtest):

do you have any idea what to do next @MATTW20 ?

OpenStudy (anonymous):

how did you get to dt=sqrt2 cos theta

OpenStudy (turingtest):

\[\frac t{\sqrt2}=\sin\theta\\t=\sqrt2\sin\theta\\dt=\sqrt2\cos\theta d\theta\]

OpenStudy (anonymous):

i see. needed that

OpenStudy (anonymous):

now you substitute

OpenStudy (turingtest):

right, and don't forget to sub for t^2

OpenStudy (anonymous):

what am i substituting in for there?

OpenStudy (turingtest):

I gave you a formula for t in the above 3 lines, you should be able to get t^2 from that

OpenStudy (anonymous):

oh i see

OpenStudy (anonymous):

so 2sin^2 theta

OpenStudy (turingtest):

oh for t^2, yes

OpenStudy (anonymous):

so then\[\frac{ \sqrt{2}\cos \theta }{ 2\sin ^{2}\theta \sqrt{1-\sin^2\theta } }\]

OpenStudy (turingtest):

yes, and don't forget we still have a factor of \[\frac1{\sqrt2}\]outside the integral, so that cancels with the one on top

OpenStudy (anonymous):

ok

OpenStudy (turingtest):

simplify as much as possible, what do you get?

OpenStudy (anonymous):

1/2sin^2 theta?

OpenStudy (turingtest):

right, integral of which is...?

OpenStudy (turingtest):

might want to use one more trig identity before you integrate...

OpenStudy (anonymous):

take out t^2 common out of root and substitute 1/t^2=c

OpenStudy (anonymous):

-1/2tan theta

OpenStudy (turingtest):

cot...

OpenStudy (anonymous):

right

OpenStudy (turingtest):

do you know how to get back to terms of t?

OpenStudy (anonymous):

this is going to sound really bad but no

OpenStudy (turingtest):

no, it's one of the more bizarre parts of trig subs if you ask me... you have to find the triangle that corresponds to your initial substitution

OpenStudy (turingtest):

so int this case our initial sub was\[\frac t{\sqrt2}=\sin\theta\]from which we can draw a triangle

OpenStudy (anonymous):

oh i have that actually 

OpenStudy (turingtest):

right, so for you, u=t, and a=?

OpenStudy (anonymous):

sqrt 2

OpenStudy (turingtest):

yup, so cot(theta)=?

OpenStudy (anonymous):

sqrt(2-t^2)/t^2

OpenStudy (turingtest):

great, and so our final solution is...?

OpenStudy (anonymous):

well what are we doing now?

OpenStudy (turingtest):

putting back the -1/2 factor that we have out in front and putting a +C is all

OpenStudy (anonymous):

oh yeah i forgot about that -1/2

OpenStudy (anonymous):

alright cool

OpenStudy (turingtest):

cool, hope you followed that well enough to tackle the rest on your own :)

OpenStudy (anonymous):

that was actually a pretty complicated problem imo

OpenStudy (turingtest):

trig subs seem really complex at first, but once you get used to them they are pretty formulaic

OpenStudy (anonymous):

ummmmm what should i look for next time because i didn't think it was that obvious it was a trig substitution problem

OpenStudy (turingtest):

when you have \[\frac1{\sqrt{a^2\pm x^2}}\] trig sub should come to mind that said, there may have been a really smart sub we could have used to get the same answer in an easier way, but this is what came to me by parts would not work though, and finding the right u-sub to make an integral like this easy is often as tricky as just going ahead with the trig sub, so it's a matter of what you are more comfortable with and stuff

OpenStudy (anonymous):

ok thnx a lot

OpenStudy (turingtest):

welcome!